OpenStudy (anonymous):

the point (-4,-7) is on the terminal side of angle theta, find the cos theta and cot theta

OpenStudy (ghazi):

 hypotenuse= \[\sqrt{(-7)^2+(-4)^2}=\sqrt{49+16}=\sqrt{65}\] so \[\cos \theta = \frac{ -4 }{ \sqrt{65} }\]

OpenStudy (ghazi):

\[\cot \theta =\frac{ -4 }{ -7 }=\frac{ 4 }{ 7 }\]

OpenStudy (anonymous):

ok thanks

OpenStudy (ghazi):

:D